OpenStudy (anonymous):

Assume that y varies inversely as x. If x = 12 when y = 9, what is x when y = -3? 108 -36 27 -4

Directrix (directrix):

-36 ==== In an inverse variation, the values of the two variables change in an opposite manner - as one value increases, the other decreases. x*y = 12*9 = 108 x (-3) = 108 x = -36